OpenStudy (fibonaccichick666):

Just need a quick check. ODE \[y"+16y'+320=0\] I'm getting \[r=-8+4i~,~-8-4i\] Is that right?

OpenStudy (fibonaccichick666):

@hartnn do you mind?

OpenStudy (fibonaccichick666):

@dan815 could you?

OpenStudy (dan815):

y=e^rx y'=re^rx y''=r^2*e^rx y''+16y'+320=0 e^rx(r^2+16r+320)=0 (r+8)^2-64+320=0 r=+/-sqrt-256 -8

OpenStudy (dan815):

r=-2^(9/2) i -8 r=2^(9/2)-8

OpenStudy (fibonaccichick666):

I see my mistake... thank you

OpenStudy (dan815):

ok np :)

OpenStudy (fibonaccichick666):

I more question in a second if you don't mind

OpenStudy (dan815):

okk

OpenStudy (fibonaccichick666):

so r=-8+-16i

OpenStudy (fibonaccichick666):

lol why?

OpenStudy (dan815):

right! it was 8/2

OpenStudy (dan815):

my bbaadd

OpenStudy (fibonaccichick666):

oh, I saw that I thought 4 times 320 was 320 that was my error lol

OpenStudy (dan815):

i forgot my middle finger is 8 not 9, my finger counting skills have begun to fail me

OpenStudy (fibonaccichick666):

lol as have all of ours

OpenStudy (fibonaccichick666):

ok so if I can just check my polar forms

OpenStudy (dan815):

u wanna write in re^ix form?

OpenStudy (fibonaccichick666):

yep for the general sln I have to

OpenStudy (dan815):

okk

OpenStudy (dan815):

r=mag of the vector <8,4>

OpenStudy (dan815):

and thet = tan^-1(4/-8)

OpenStudy (fibonaccichick666):

radius is 5 sqrt10

OpenStudy (dan815):

that looks big

OpenStudy (fibonaccichick666):

how'd you get 4?

OpenStudy (fibonaccichick666):

I like that better

OpenStudy (dan815):

oh my bad i was lookign at ur top solution on post

OpenStudy (dan815):

in that case 5root10 looks about right!

OpenStudy (fibonaccichick666):

still I'm getting arctan(-2)...???

OpenStudy (dan815):

yeah thats right

OpenStudy (dan815):

whats wrong with that

OpenStudy (fibonaccichick666):

but I don't know the answer to that

OpenStudy (dan815):

oh lol

OpenStudy (dan815):

u cant use calculator??

OpenStudy (fibonaccichick666):

no no calculators allowed

OpenStudy (dan815):

well thats not nice umm

OpenStudy (fibonaccichick666):

ughhh did I do something wrong setting it up then? hmm

OpenStudy (dan815):

when is cos theta half of sin tehta then?

OpenStudy (fibonaccichick666):

I guess can I give you the whole problem to look at?

OpenStudy (dan815):

ok

OpenStudy (fibonaccichick666):

A mass weighing 2lb stretches a spring 1.2in. The mass is pushed 6in. downward and released. The motion takes place in a medium with damping constant 1lb s/ft. Find the position of the mass at any time t\(\ge\)0

OpenStudy (fibonaccichick666):

so w=2lb m=2/32=1/16 k=w/x=2/.1=20 x=1.2in=.1ft \(\gamma\)=1 lb s/ft

OpenStudy (fibonaccichick666):

my eq should be of the form: \[my"+\gamma y'+ky=g(t)\]

OpenStudy (fibonaccichick666):

so i set it up: \[\frac{1}{16}y"+y'+20y=0\] which becomes \[y''+16y'+320y=0\]

OpenStudy (dan815):

re u sure k=w/x

OpenStudy (dan815):

is that some thing arbritray they gave u

OpenStudy (fibonaccichick666):

no hookes law mg=kL

OpenStudy (dan815):

oh really

OpenStudy (dan815):

wait what

OpenStudy (fibonaccichick666):

that's what the book said...

OpenStudy (dan815):

wait do u mean F=-kx

OpenStudy (fibonaccichick666):

mg − kL = 0.

OpenStudy (fibonaccichick666):

according to the book

OpenStudy (dan815):

by hookes law u mean like the acceleration relationship wrt to postion right

OpenStudy (fibonaccichick666):

yea

OpenStudy (fibonaccichick666):

is it supposed to be negative?

OpenStudy (dan815):

well its usually written in negative unless they refine L

OpenStudy (fibonaccichick666):

I mean is k supposed to be negative in the end?

OpenStudy (dan815):

but u cannot get K from that eqution though

OpenStudy (dan815):

k has to be suppplied

OpenStudy (fibonaccichick666):

we have to find k

OpenStudy (dan815):

oh hoho

OpenStudy (dan815):

well then!

OpenStudy (dan815):

m dy^2/dt= -k*(Y_o-y) + D*(dy/dt)

OpenStudy (dan815):

Y_o is equlibrium pos

OpenStudy (fibonaccichick666):

wha???

OpenStudy (dan815):

that is your normal equation with D for damping

OpenStudy (fibonaccichick666):

I've never seen that

OpenStudy (dan815):

see how the damping is taking away from -kdeltaposition that is gonna slow down our accerlation

OpenStudy (dan815):

m dy^2/dt= -k*(Y_o-y) this is F=-kx

OpenStudy (dan815):

i just added damping term to it

OpenStudy (fibonaccichick666):

but we aren't supposed to?

OpenStudy (dan815):

fibonaccci!

OpenStudy (dan815):

ill hit u -.-

OpenStudy (fibonaccichick666):

I''m sorry, we never covered that in class!!! I'm a wee bit confused

OpenStudy (fibonaccichick666):

we only covered mg=kx

OpenStudy (dan815):

okay fine lets get something straight here hookes law

OpenStudy (dan815):

is ma=kdeltax isnt it?

OpenStudy (dan815):

are talking about the samething

OpenStudy (fibonaccichick666):

http://people.math.umass.edu/~tanguay/331/331spr10hw17sol.pdf They say it the same way here

OpenStudy (fibonaccichick666):

and yes, the elongation of the spring

OpenStudy (dan815):

how can u say mg=kx firstly its k*deltaX and the whole point is delta x is changing wrt to the acceration is experiences unless the k constant is the particular enlogation L when it experiences a acceraltion = gravity

OpenStudy (fibonaccichick666):

but that's what it tells me to do for spring constant at equilibrium

OpenStudy (dan815):

ok ok i think u are right !! baha

OpenStudy (fibonaccichick666):

it's how they do it in the book

OpenStudy (fibonaccichick666):

lol

OpenStudy (fibonaccichick666):

god... I'm correcting my test so I can go beg for points at like 9a

OpenStudy (fibonaccichick666):

but I keep screwing something up

OpenStudy (fibonaccichick666):

grrrrrr

OpenStudy (dan815):

i see ur mistake i think

OpenStudy (fibonaccichick666):

what? please tell me

OpenStudy (dan815):

mg=kL 2*9.81/1.2=k

OpenStudy (dan815):

wait thats in lb

OpenStudy (dan815):

whats the gravity thing in lb?

OpenStudy (fibonaccichick666):

32 not 9.81

OpenStudy (dan815):

okay

OpenStudy (fibonaccichick666):

and change that 1.2in to feet

OpenStudy (fibonaccichick666):

where'd the 2 come from?

OpenStudy (dan815):

u said 1.2in.